# Quantum many-body scars with unconventional superconducting pairing symmetries via multi-body interactions

### v2

@@ VI Generalizations
In this section, we present several generalizations of our approach to construct superconducting scar states with the multi-body interactions: generalized off-site $\eta$-pairing states (Sec. VI.1), $f$-wave $\eta$-pairing states on the honeycomb lattice (Sec. VI.2), Yang’s $\eta$-pairing states in a nearest-neighbor interacting system (Sec. VI.3), and spinless $s$-wave $\eta$-pairing states in the one-dimensional system (Sec. VI.4). Furthermore, we can also extend the form of the model Hamiltonian that has the unconventional $\eta$-pairing eigenstates. In Sec. VI.5, we show that the $d$-wave $\eta$-pairing state also becomes an eigenstate in a model with higher multi-body interaction. In Sec. VI.6, we argue that the eigenenergy of the unconventional $\eta$-pairing states can be controlled by an effective chemical potential in terms of pair numbers. In Sec. VI.7, we discuss the case of the open boundary condition, for which we determine the condition that the unconventional $\eta$-pairing states remain to be the eigenstates.
@@ VI.1 General pairing symmetry and long-range pairing
The extended Hubbard model with the multi-body interactions in Eq. (29) has not only the $d$-wave $\eta$-pairing state but also several other $\eta$-pairing states with different pairing symmetries as energy eigenstates, since the energy cancellation holds for each Fock basis. Let us consider $s$-, $p_{x}$-, and $p_{y}$-wave pairing symmetries, for which $\eta$-pairing operators are defined by
@@
where $\nu_{\gamma}$ ($\gamma=s,p_{x},p_{y},d$) denotes the number of pairs with the $\gamma$-wave pairing symmetry. These $\eta$-pairing states are eigenstates of $\mathcal{H}_{\mathrm{extH}}$ in Eq. (29), which are $\binom{N/2+3}{3}$-fold degenerate with zero eigenenergy. Note that an arbitrary linear combination of the general $\eta$-pairing states [Eq. (36)] is also an eigenstate of $\mathcal{H}_{\mathrm{extH}}$ in Eq. (29) (including, e.g., the $p_{x}+ip_{y}$-wave state and the $s+d$-wave state), since they are all degenerate at zero energy. In order to induce the $d$-wave $\eta$-pairing state by a certain excitation protocol, it will be better if the $d$-wave pairing state is energetically separated from the other degenerate states. In this paper, we leave it an open issue how to break the degeneracy.
While we have focused on the two-dimensional square lattice so far, most of the results can be straightforwardly extended to other dimensions and other bipartite lattice systems. The bipartite lattice condition is necessary for the $\eta$-pairing states to be eigenstates of the kinetic term $\mathcal{H}_{\mathrm{kin}}$ [Eq. (21)]. We can define the interaction term $\mathcal{H}_{\mathrm{int}}$ [Eq. (30)] and the neighboring particle existence operator $\mathbb{n}_{i\sigma}$ in Eqs. (25) and (26) irrespective of dimensions and lattice structures. Therefore, one can use the Hamiltonian of Eq. (29), for example, on the one-dimensional chain, honeycomb, body-centered-cubic, and diamond lattices.
For the one-dimensional chain, we can obtain the spin-singlet $p$-wave pairing state $(\eta_{p_{x}}^{+})^{N/2}|0\rangle$ with $\eta_{p_{x}}^{+}$ defined in Eq. (34), as a QMBS state. The scarred Hamiltonian consists of interaction terms involving up to a four-body interaction,
@@
To make the long-range $\eta$-pairing states exact eigenstates of the Hamiltonian, one should replace the form of the multi-body interaction in Eq. (28) by $-U\sum_{\sigma}n_{i\uparrow}n_{i\downarrow}\mathbb{n}_{i\sigma}^{(n)}/2$ with
@@
In this way, the $\eta$-pairing states, which consist of the long-range pairs and have different types of parity and spin symmetries, also become energy eigenstates due to the longer-range multi-body interaction. It should be noted that our method of realizing superconducting scar states using multi-body interactions is only applicable to pairs of particles separated by a fixed distance.
@@ VI.2 $f$-wave $\eta$-pairing state on the honeycomb lattice
Figure 8: Generalized $\eta$-pairing states on the honeycomb lattice. (a) Six kinds of next-nearest-neighbor pairs (connected by the dash-dotted lines) in the $\eta$-pairing states. The white (grey) circles represent $\mathrm{A}$ ($\mathrm{B}$) sublattice sites. $\bm{\tau}_{\gamma}$ and $\bm{R}_{\gamma}$ ($\gamma=1,2,3$) denote the bond and lattice vectors, respectively. A pair on the $\mathrm{B}$ sublattice (connected by the blue dash-dotted line) has the opposite sign of the wavefunction with respect to those on the $\mathrm{A}$ sublattice. (b) Energy cancellation between the two-body (black lines) and multi-body (filled areas) interactions acting on overlapped pairs.
Let us consider the multi-body interacting model defined by Eq. (29) on the honeycomb lattice, where an $\eta$-pairing state with the $f$-wave pairing symmetry becomes an eigenstate [see Fig. 1(c)]. The lattice structure is defined as follows. We label sublattices by $\mathrm{A}$ and $\mathrm{B}$, as shown in Fig. 8(a). Let $\bm{\tau}_{\gamma}$ ($\gamma=1,2,3$) be the bond vectors and $\bm{R}_{\gamma}$ the lattice vectors, which are explicitly given by
@@
Given $g(\bm{k})=g^{*}(-\bm{k})$, the first and fourth terms and the second and third terms cancel each other, indicating that $[\mathcal{H}_{\mathrm{kin}},\eta_{\bm{\alpha}}^{+}]=0$. This is also confirmed from the commutation relation between Eqs. (53) and (62) as $\left[\mathcal{H}_{\mathrm{kint}},\eta_{\bm{\alpha}}^{+}\right]=\sum_{\bm{k}}\exp(\mathrm{i}\bm{k}\cdot\bm{\alpha})\,(|g(\bm{k})|-|g(-\bm{k})|)({c_{\bm{k},\uparrow}^{(+)}}^{\dagger}{c_{-\bm{k},\downarrow}^{(-)}}^{\dagger}+{c_{\bm{k},\uparrow}^{(-)}}^{\dagger}{c_{-\bm{k},\downarrow}^{(+)}}^{\dagger})=0$. Therefore, the $f$-wave $\eta$-pairing state is a zero-energy eigenstate of the kinetic term in Eq. (21).
By applying the neighboring particle existence operator $\mathbb{n}_{i\sigma}(2)$ defined in Eq. (39) to the case of $\mathrm{nn}(2,i)$, the $f$-wave $\eta$-pairing state also becomes an energy eigenstate of the Hamiltonian in Eq. (29). The energy increase due to the two-body interaction $U$ is canceled by the multi-body interactions between a doublon and a next-nearest neighboring particle, as shown in Fig. 8(b). Therefore, using the multi-body interactions, we can realize the $\eta$-pairing states with unconventional pairing symmetries in various lattice systems.
@@ VI.3 Yang’s $s$-wave $\eta$-pairing state in a nearest-neighbor interacting system
Yang’s $\eta$-pairing state is no longer an energy eigenstate when one adds nearest-neighbor two-body interactions to the Hubbard model in Eq. (20) due to the breaking of $\eta$-$\mathrm{SU}(2)$ symmetry [51, 52]. However, one can make Yang’s $\eta$-pairing state the energy eigenstate in nearest-neighbor interacting systems by utilizing appropriate multi-body interactions like those discussed above. The following discussions in this subsection are applicable to systems in arbitrary dimensions and with both the periodic and open boundary conditions.
@@
Therefore, utilizing the multi-body interaction, one can make Yang’s $\eta$-pairing state an energy eigenstate even in the presence of the nearest-neighbor interaction of Eq. (66).
We numerically confirm that Yang’s $\eta$-pairing state can be regarded as a QMBS state in the model with the Hamiltonian $\mathcal{H}_{\mathrm{H}}+\mathcal{H}_{\mathrm{mod}V}$ in Eq. (67). For the half-filled one-dimensional system with the periodic boundary condition of length $L$, we numerically obtain the level-spacing distribution $P(s)$ [Fig. 9(c)] in the subspace with even parity, even time-reversal symmetry, and zero total momentum. The level-spacing distribution agrees well with the Winger–Dyson distribution $P_{\mathrm{WD}}(s)$, implying that the model ($\mathcal{H}_{\mathrm{H}}+\mathcal{H}_{\mathrm{mod}V}$) is nonintegrable. We also consider the corresponding model with the open boundary condition, for which the bipartite entanglement entropy for each energy eigenstate is shown in Fig. 9(d). We find that Yang’s $\eta$-pairing state, indicated by the red circle, has much smaller entanglement entropy than the other thermal states (with $S_{\mathrm{A}}\gtrsim 3\ln 2$). Therefore, we conclude that Yang’s $\eta$-pairing state is a nonthermal energy eigenstate (i.e., a QMBS state) in the multi-body interacting system with the Hamiltonian $\mathcal{H}_{\mathrm{H}}+\mathcal{H}_{\mathrm{mod}V}$.
@@ VI.4 The spinless $s$-wave $\eta$-pairing state in the one-dimensional system
Figure 10: (a) Nearest-neighbor level-spacing distribution $P(s)$ for the one-dimensional spinless Hubbard model with the three-body interaction [Eq. (13)] with $L=20$. The blue and red curves correspond to the Poisson and Wigner-Dyson distribution, respectively. (b) Entanglement entropy spectrum of the energy eigenstates in the corresponding model with $L=16$. The red circle indicates the spinless $s$-wave $\eta$-pairing state [Eq. (9)]. The parameters are set to be $t=1$, $V=2$, $W=-V/2$, $N=L/2$, and $\sigma_{\mathrm{Gauss}}=0.5$.
@@
The spinless $s$-wave $\eta$-pairing state in Eq. (9) discussed in Sec. III.1 is one of the simplest examples that show unconventional pairing symmetries, since the conventional pairings have the spin-singlet and $s$-wave symmetry. Let us first check that the spinless $s$-wave $\eta$-pairing state satisfies the conditions for a QMBS state in the extended spinless Hubbard model with the three-body interaction in Eq. (13). Figure 10(a) shows the level-spacing distribution $P(s)$ of the energy eigenstates on the half-filled periodic lattice in the subspace with odd parity and total momentum $\pi$. We observe good agreement between $P(s)$ and the Winger–Dyson distribution. Figure 10(b) shows the bipartite entanglement entropy spectrum in the open boundary chain at half filling. We find that the spinless $s$-wave $\eta$-pairing state, indicated by the red circle, has much smaller entanglement entropy than those of the other eigenstates that form a convex upward curve. Therefore, the spinless $s$-wave $\eta$-pairing state can be regarded as a QMBS state in the multi-body interacting system.
@@
Next, we clarify the relationship between the spinless Hamiltonian in Eq. (13) and our construction of scarred spinful Hamiltonians using the operator $\mathbb{n}_{i\sigma}$ in Eq. (25) introduced in Sec. III.4. The spinless system can be described by a spinful model in which all the spins of the particles are polarized to the same direction (say, $\sigma$). The spinless $\eta$-pairing operator $\eta_{\mathrm{sl}}^{+}$ in Eq. (8) corresponds to the general $\eta$-pairing operator in Eq. (1) where two fermions have the same $\sigma$-spin, i.e., $\eta_{\rm sl}^{+}=\eta_{+1,\sigma\sigma}^{+}$. The two-body interaction term $\mathcal{H}_{V}$ in Eq. (11) corresponds to a nearest-neighbor interaction in a spinful model, $\sum_{i}n_{i\sigma}n_{i+1,\sigma}$. The scarred Hamiltonian including the three-body interaction in Eq. (15) can be rewritten as $V\sum_{i}n_{i\sigma}\mathbb{n}_{i\sigma}/2$ using the neighboring particle existence operator $\mathbb{n}_{i\sigma}$ in the one-dimensional chain, which is explicitly represented by $\mathbb{n}_{i\sigma}=1-(1-n_{i-1,\sigma})(1-n_{i+1,\sigma})=n_{i-1,\sigma}+n_{i+1,\sigma}-n_{i-1,\sigma}n_{i+1,\sigma}$. The term $V\sum_{i}n_{i\sigma}\mathbb{n}_{i\sigma}/2$ means that each fermion pair always has the constant energy $V$ in any pair configuration, no matter how close the pairs are coming to each other. We will discuss the spinful version of this term in more detail in Sec. VI.6.
VI.5 Higher multi-body interactions

### v3

@@ VI Generalizations
In this section, we present several generalizations of our approach to construct superconducting scar states with the multibody interactions: generalized off-site $\eta$-pairing states (Sec. VI.1), $f$-wave $\eta$-pairing states on the honeycomb lattice (Sec. VI.2), Yang’s $\eta$-pairing states in a nearest-neighbor interacting system (Sec. VI.3), and spinless $s$-wave $\eta$-pairing states in the one-dimensional system (Sec. VI.4). Furthermore, we can also extend the form of the model Hamiltonian that has the unconventional $\eta$-pairing eigenstates. In Sec. VI.5, we show that the $d$-wave $\eta$-pairing state also becomes an eigenstate in a model with higher multibody interaction. In Sec. VI.6, we argue that the eigenenergy of the unconventional $\eta$-pairing states can be controlled by an effective chemical potential in terms of pair numbers. In Sec. VI.7, we discuss the case of the open boundary condition, for which we determine the condition that the unconventional $\eta$-pairing states remain to be the eigenstates.
@@ VI.1 General pairing symmetry and long-range pairing
The extended Hubbard model with the multibody interactions in Eq. (29) has not only the $d$-wave $\eta$-pairing state but also several other $\eta$-pairing states with different pairing symmetries as energy eigenstates, since the energy cancellation holds for each Fock basis. Let us consider $s$-, $p_{x}$-, and $p_{y}$-wave pairing symmetries, for which $\eta$-pairing operators are defined by
@@
where $\nu_{\gamma}$ ($\gamma=s,p_{x},p_{y},d$) denotes the number of pairs with the $\gamma$-wave pairing symmetry. These $\eta$-pairing states are eigenstates of $\mathcal{H}_{\mathrm{extH}}$ in Eq. (29), which are $\binom{N/2+3}{3}$-fold degenerate with zero eigenenergy. Note that an arbitrary linear combination of the general $\eta$-pairing states [Eq. (36)] is also an eigenstate of $\mathcal{H}_{\mathrm{extH}}$ in Eq. (29) (including, e.g., the $p_{x}+ip_{y}$-wave state and the $s+d$-wave state), since they are all degenerate at zero energy. In order to induce the $d$-wave $\eta$-pairing state by a certain excitation protocol, it will be better if the $d$-wave pairing state is energetically separated from the other degenerate states. In this paper, we leave it an open issue how to break the degeneracy, which would be effectively addressed by numerical methods for systematically constructing scarred Hamiltonians [84, 85].
While we have focused on the two-dimensional square lattice so far, most of the results can be straightforwardly extended to other dimensions and other bipartite lattice systems. The bipartite lattice condition is necessary for the $\eta$-pairing states to be eigenstates of the kinetic term $\mathcal{H}_{\mathrm{kin}}$ [Eq. (21)]. We can define the interaction term $\mathcal{H}_{\mathrm{int}}$ [Eq. (30)] and the neighboring particle existence operator $\mathbb{n}_{i\sigma}$ in Eqs. (25) and (26) irrespective of dimensions and lattice structures. Therefore one can use the Hamiltonian of Eq. (29), for example, on the one-dimensional chain, honeycomb, body-centered-cubic, and diamond lattices.
For the one-dimensional chain  [^1] In the one-dimensional case, the spin-singlet and parity-odd $\eta$-pairing operator is mentioned in Ref. [54]. Specifically, it is shown that the state obtained by acting with the $\eta_{p_{x}}^{+}$ operator [Eq. (34)] only once on Yang’s $\eta$-pairing state [Eq. (65)] becomes an eigenstate of the Hirsch model. To extend this result to states where the $\eta_{p_{x}}^{+}$ operator is applied multiple times or to higher-dimensional systems, the multibody interactions proposed in this study are required., we can obtain the spin-singlet $p$-wave pairing state $(\eta_{p_{x}}^{+})^{N/2}|0\rangle$ with $\eta_{p_{x}}^{+}$ defined in Eq. (34), as a QMBS state. The scarred Hamiltonian consists of interaction terms involving up to a four-body interaction,
@@
To make the long-range $\eta$-pairing states exact eigenstates of the Hamiltonian, one should replace the form of the multibody interaction in Eq. (28) by $-U\sum_{\sigma}n_{i\uparrow}n_{i\downarrow}\mathbb{n}_{i\sigma}^{(n)}/2$ with
@@
In this way, the $\eta$-pairing states, which consist of the long-range pairs and have different types of parity and spin symmetries, also become energy eigenstates due to the longer-range multibody interaction. It should be noted that our method of realizing superconducting scar states using multibody interactions is only applicable to pairs of particles separated by a fixed distance.
@@ VI.2 $f$-wave $\eta$-pairing state on the honeycomb lattice
Figure 8: Generalized $\eta$-pairing states on the honeycomb lattice. (a) Six kinds of next-nearest-neighbor pairs (connected by the dash-dotted lines) in the $\eta$-pairing states. The white (grey) circles represent $\mathrm{A}$ ($\mathrm{B}$) sublattice sites. $\bm{\tau}_{\gamma}$ and $\bm{R}_{\gamma}$ ($\gamma=1,2,3$) denote the bond and lattice vectors, respectively. A pair on the $\mathrm{B}$ sublattice (connected by the blue dash-dotted line) has the opposite sign of the wavefunction with respect to those on the $\mathrm{A}$ sublattice. (b) Energy cancellation between the two-body (black lines) and multibody (filled areas) interactions acting on overlapped pairs.
Let us consider the multibody interacting model defined by Eq. (29) on the honeycomb lattice, where an $\eta$-pairing state with the $f$-wave pairing symmetry becomes an eigenstate [see Fig. 1(c)]. The lattice structure is defined as follows. We label sublattices by $\mathrm{A}$ and $\mathrm{B}$, as shown in Fig. 8(a). Let $\bm{\tau}_{\gamma}$ ($\gamma=1,2,3$) be the bond vectors and $\bm{R}_{\gamma}$ the lattice vectors, which are explicitly given by
@@
Given $g(\bm{k})=g^{*}(-\bm{k})$, the first and fourth terms and the second and third terms cancel each other, indicating that $[\mathcal{H}_{\mathrm{kin}},\eta_{\bm{\alpha}}^{+}]=0$. This is also confirmed from the commutation relation between Eqs. (53) and (62) as $\left[\mathcal{H}_{\mathrm{kin}},\eta_{\bm{\alpha}}^{+}\right]=\sum_{\bm{k}}\exp(\mathrm{i}\bm{k}\cdot\bm{\alpha})\,(|g(\bm{k})|-|g(-\bm{k})|)({c_{\bm{k},\uparrow}^{(+)}}^{\dagger}{c_{-\bm{k},\downarrow}^{(-)}}^{\dagger}+{c_{\bm{k},\uparrow}^{(-)}}^{\dagger}{c_{-\bm{k},\downarrow}^{(+)}}^{\dagger})=0$. Therefore the $f$-wave $\eta$-pairing state is a zero-energy eigenstate of the kinetic term in Eq. (21).
By applying the neighboring particle existence operator $\mathbb{n}_{i\sigma}(2)$ defined in Eq. (39) to the case of $\mathrm{nn}(2,i)$, the $f$-wave $\eta$-pairing state also becomes an energy eigenstate of the Hamiltonian in Eq. (29). The energy increase due to the two-body interaction $U$ is canceled by the multibody interactions between a doublon and a next-nearest neighboring particle, as shown in Fig. 8(b). Therefore, using the multibody interactions, we can realize the $\eta$-pairing states with unconventional pairing symmetries in various lattice systems.
@@ VI.3 Yang’s $s$-wave $\eta$-pairing state in a nearest-neighbor interacting system
Yang’s $\eta$-pairing state is no longer an energy eigenstate when one adds nearest-neighbor two-body interactions to the Hubbard model in Eq. (20) due to the breaking of $\eta$-$\mathrm{SU}(2)$ symmetry [51, 52]. However, one can make Yang’s $\eta$-pairing state the energy eigenstate in nearest-neighbor interacting systems by utilizing appropriate multibody interactions like those discussed above. In Ref. [54], it has been studied how to make Yang’s $\eta$-pairing state a QMBS state, where the multibody interaction Hamiltonian presented below can be reproduced. Here, we include this example to show the role of the multibody interaction discussed above. The following discussions in this subsection are applicable to systems in arbitrary dimensions and with both the periodic and open boundary conditions.
@@
Therefore, utilizing the multibody interaction, one can make Yang’s $\eta$-pairing state an energy eigenstate even in the presence of the nearest-neighbor interaction of Eq. (66).
We numerically confirm that Yang’s $\eta$-pairing state can be regarded as a QMBS state in the model with the Hamiltonian $\mathcal{H}_{\mathrm{H}}+\mathcal{H}_{\mathrm{mod}V}$ in Eq. (67). For the half-filled one-dimensional system with the periodic boundary condition of length $L$, we numerically obtain the level-spacing distribution $P(s)$ [Fig. 9(c)] in the subspace with even parity, even time-reversal symmetry, and zero total momentum. The level-spacing distribution agrees well with the Winger–Dyson distribution $P_{\mathrm{WD}}(s)$, implying that the model ($\mathcal{H}_{\mathrm{H}}+\mathcal{H}_{\mathrm{mod}V}$) is nonintegrable. We also consider the corresponding model with the open boundary condition, for which the bipartite entanglement entropy for each energy eigenstate is shown in Fig. 9(d). We find that Yang’s $\eta$-pairing state, indicated by the red circle, has much smaller entanglement entropy than the other thermal states (with $S_{\mathrm{A}}\gtrsim 3\ln 2$). Therefore we conclude that Yang’s $\eta$-pairing state is a nonthermal energy eigenstate (i.e., a QMBS state) in the multibody interacting system with the Hamiltonian $\mathcal{H}_{\mathrm{H}}+\mathcal{H}_{\mathrm{mod}V}$.
@@ VI.4 The spinless $s$-wave $\eta$-pairing state in the one-dimensional system
Figure 10: (a) Nearest-neighbor level-spacing distribution $P(s)$ for the one-dimensional spinless Hubbard model with the three-body interaction [Eq. (13)] with $L=20$. The blue and red curves correspond to the Poisson and Wigner–Dyson distribution, respectively. (b) Entanglement entropy spectrum of the energy eigenstates in the corresponding model with $L=16$. The red circle indicates the spinless $s$-wave $\eta$-pairing state [Eq. (9)]. The parameters are set to be $t=1$, $V=2$, $W=-V/2$, $N=L/2$, and $\sigma_{\mathrm{Gauss}}=0.5$.
@@
The spinless $s$-wave $\eta$-pairing state in Eq. (9) discussed in Sec. III.1 is one of the simplest examples that show unconventional pairing symmetries, since the conventional pairings have the spin-singlet and $s$-wave symmetry. Let us first check that the spinless $s$-wave $\eta$-pairing state satisfies the conditions for a QMBS state in the extended spinless Hubbard model with the three-body interaction in Eq. (13). Figure 10(a) shows the level-spacing distribution $P(s)$ of the energy eigenstates on the half-filled periodic lattice in the subspace with even parity and total momentum $\pi$. We observe good agreement between $P(s)$ and the Winger–Dyson distribution. Figure 10(b) shows the bipartite entanglement entropy spectrum in the open boundary chain at half filling. We find that the spinless $s$-wave $\eta$-pairing state, indicated by the red circle, has much smaller entanglement entropy than those of the other eigenstates that form a convex upward curve. Therefore the spinless $s$-wave $\eta$-pairing state can be regarded as a QMBS state in the multibody interacting system.
@@
Next, we clarify the relationship between the spinless Hamiltonian in Eq. (13) and our construction of scarred spinful Hamiltonians using the operator $\mathbb{n}_{i\sigma}$ in Eq. (25) introduced in Sec. III.4. The spinless system can be described by a spinful model in which all the spins of the particles are polarized to the same direction (say, $\sigma$). The spinless $\eta$-pairing operator $\eta_{\mathrm{sl}}^{+}$ in Eq. (8) corresponds to the general $\eta$-pairing operator in Eq. (1) where two fermions have the same $\sigma$-spin, i.e., $\eta_{\mathrm{sl}}^{+}=\eta_{+1,\sigma\sigma}^{+}$. The two-body interaction term $\mathcal{H}_{V}$ in Eq. (11) corresponds to a nearest-neighbor interaction in a spinful model, $\sum_{i}n_{i\sigma}n_{i+1,\sigma}$. The scarred Hamiltonian including the three-body interaction in Eq. (15) can be rewritten as $V\sum_{i}n_{i\sigma}\mathbb{n}_{i\sigma}/2$ using the neighboring particle existence operator $\mathbb{n}_{i\sigma}$ in the one-dimensional chain, which is explicitly represented by $\mathbb{n}_{i\sigma}=1-(1-n_{i-1,\sigma})(1-n_{i+1,\sigma})=n_{i-1,\sigma}+n_{i+1,\sigma}-n_{i-1,\sigma}n_{i+1,\sigma}$. The term $V\sum_{i}n_{i\sigma}\mathbb{n}_{i\sigma}/2$ means that each fermion pair always has the constant energy $V$ in any pair configuration, no matter how close the pairs are coming to each other. We will discuss the spinful version of this term in more detail in Sec. VI.6.
VI.5 Higher multibody interactions